# Blowin’ in the non-isothermal wind: core-powered mass loss with hydrodynamic radiative transfer

### v1

@@ 3.1 Sensitivity of mass loss rates to opacity
Figure 2: Temperature (top) and density (bottom) profiles as functions of radius, in core radii $R_{\mathrm{c}}$, and corresponding mass loss rates (right) for seven values of the opacity ratio ranging from 0.03 to 30 (different colors from yellow to purple). All profiles have a fixed radiative-convective boundary location $R_{\mathrm{rcb}}=2R_{\mathrm{c}}=1.91\times 10^{9}$ cm, planet mass $M_{\mathrm{p}}=5M_{\oplus}$, atmospheric mass fraction $f=0.025$, and incident flux equivalent to an equilibrium temperature $T_{\mathrm{eq}}=1000\,$K (shown by the gray horizontal line on the top panel). T-shaped markers represent the transit radius, while squares represent the sonic radius. Shown in red dashed lines are the equivalent profiles and mass loss rates for an isothermal radiative region at the analytic interior temperature, $T_{\mathrm{in}}=T_{\mathrm{eq}}/2^{1/4}$. Increasing the opacity ratio leads to different temperature and density profiles and therefore to mass loss rates that vary two orders of magnitude over plausible values. The mass loss rate for the $\gamma=1$ profile are similar to those for an isothermal outer atmosphere at $T_{\mathrm{in}}$.
@@
In Fig. 2, we compare the thermal (top left) and density (bottom left) profiles and mass loss rates (right) as functions of the ratio of the opacity to incident visible and thermal radiation, $\gamma$. Each color is a run with a value of this ratio ranging from 0.3 to 30, with $\gamma$ increasing as the colors get darker. All seven runs have a fixed radiative-convective boundary radius, $R_{\mathrm{rcb}}=2R_{\mathrm{c}}=1.91\times 10^{9}$ cm, a fixed initial atmospheric mass of $0.025M_{\mathrm{c}}$, a planet mass $M_{\mathrm{p}}=5M_{\oplus}$, and a fixed incident flux equivalent to an equilibrium temperature of 1000 K, which is shown by the gray horizontal line. For each $\gamma$ value, the solid line represents the output of the aiolos code, while the dotted line represents the semi-analytic extrapolation into the optically thick region described in Section 2.6. T-shaped markers represent the transit radius, $R_{\mathrm{trans}}$ (Eq. 8), while squares represent the sonic radius, $R_{\mathrm{s}}$ (Eq. 11). The red dashed lines and red diamond represent the hydrostatic profile of an isothermal outer atmosphere with the same radiative-convective boundary radius and atmospheric mass and an constant temperature equal to $T_{\mathrm{in}}=T_{\mathrm{eq}}/2^{1/4}$.
@@
In general, the numerical thermal profiles reproduce the characteristics described in Section 2, thus looking broadly similar to the schematic in Fig. 1. As the opacity ratio $\gamma$ increases, the outer region, optically thin to incident radiation, heats up, with temperatures increasing by a factor of 7 over the studied 1000-fold opacity ratio increase. The outer temperatures range from well below the equilibrium temperature to well above it, and are well-predicted by the $T_{\mathrm{out}}=T_{\mathrm{eq}}(\gamma/4)^{1/4}$ relation of Guillot (2010); Schulik & Booth (2023). Meanwhile the inner region, optically thick to outgoing radiation, reaches lower temperatures as $\gamma$ is increased, decreasing by 100 K over the studied range. In all cases, the inner, optically thick radiative region is a few hundred kelvins cooler than the equilibrium temperature. This temperature is also lower by $\sim 70-200$ K than the analytic lower limit of $T_{\mathrm{in}}$ discussed in Schulik & Booth (2023). The minimum temperature also increases monotonically with $\gamma$, rather than being highest at $\gamma\sim 1$. This is because the inner temperature approximation of Parmentier & Guillot (2014) was calculated assuming plane-parallel geometry. However, in small planets the scale height is a non-negligible fraction of the planet radius. Therefore, in our simulations the different surface areas of the absorption and emission surfaces lead to a decrease in temperature as $\kappa_{P,\odot}$ increases, since the same radiation is absorbed over a progressively larger area.
@@ 4.2 Comparison to direct observations
A more direct test of the thermal structure of sub-Neptunes and its effects on planet evolution could come from atmospheric escape measurements. Signatures of escaping atmospheres have been observed directly for sub-Neptunes. Detections consistent with atmospheric escape for small exoplanets have been made using hydrogen Ly $\alpha$ 121 nm in the UV (e.g. Kulow et al., 2014; Ehrenreich et al., 2015) and He 1083 nm in the NIR (e.g. Mansfield et al., 2018) lines (see Dos Santos (2023) for a recent review). If these planets are undergoing bolometrically-driven escape, measurements of the mass loss rate will constrain the temperature of the upper atmosphere, and therefore the atmospheric opacity ratio. However, the mass loss rate is difficult to constrain precisely from these measurements, due to degeneracies between escape velocity and ionization by the stellar wind (Owen et al., 2023; Schreyer et al., 2023). But using existing models, these measurements have been fit by isothermal mass loss rate estimates for sub-Neptunes as low as $10^{9}$ g s-1 (Ehrenreich et al., 2015; Mansfield et al., 2018). This measured rate is within the range of escape rates we predict on billion year timescales for some planet-$\gamma$ combinations, indicating that directly testing upper atmospheric structure with escape observations is within reach of existing telescopic capabilities. We facilitate comparisons to our results by tabulating the mass loss rates we obtain with aiolos as functions of planet mass, equilibrium temperature, value of $\gamma$, and transit radius and including them as machine-readable Supplementary Data. We describe these tables in Appendix B.
@@ 4.3 Opacity variations in space and time
Figure 9: Temperature (top) and density (bottom) profiles as functions of radius, in core radii $R_{\mathrm{c}}$, and corresponding mass loss rates (right) for six choices of Planck opacities. In the darker colors are three profiles with the default thermal Planck opacity presented in the results section, $\kappa_{\mathrm{P,therm}}=7.5$ cm2 g-1, and solar opacities such that the ratio between the two, $\gamma$ is 0.3 (green), 3 (blue), and 30 (purple). These profiles were previously shown in Fig. 2. The lighter shades of each color correspond to runs with the same $\gamma$ values, but all opacities lowered by a factor of 10, such that $\kappa_{\mathrm{P,therm}}=0.75$ cm2 g-1. In the right panel, the mass loss rates for each particular value of $\gamma$ are linked by solid lines. All opacities listed in text on the figure are in cm2 g-1. All profiles have a fixed radiative-convective boundary location $R_{\mathrm{rcb}}=2R_{\mathrm{c}}=1.91\times 10^{9}$ cm, atmospheric mass fraction $f=0.025$, and incident flux equivalent to an equilibrium temperature $T_{\mathrm{eq}}=1000\,$K (shown by the gray horizontal line on the top panel). Shown in red dashed lines are the equivalent profiles and mass loss rates for an isothermal radiative region at the analytic interior temperature, $T_{\mathrm{in}}=T_{\mathrm{eq}}/2^{1/4}$. As in Fig. 2, T-shaped markers represent the transit radius, while squares represent the sonic radius. Decreasing the opacities moves the $\tau=1$ surfaces inward, and therefore the transitions between $T_{\mathrm{in}}$ and $T_{\mathrm{out}}$ move inward for each $\gamma$. For $\gamma>1$, where $T_{\mathrm{out}}>T_{\mathrm{in}}$, this increases the mass loss rate, up to a factor of 5 for the $\gamma=30$ case shown, while it lowers the mass loss rate for $\gamma<1$. The value of $\gamma$ still sets the outer temperatures and therefore remains the main control on the mass loss rate.
@@
In order to ascertain that our results are independent to first-order of our exact choice of the thermal Planck opacity $\kappa_{\mathrm{P,therm}}$, we run simulations with a lower value, $\kappa_{\mathrm{P,therm}}=0.75$ cm2 g-1, shown in Figure 9. Here we show the temperature and density profiles as functions of radius, as well as the mass loss rate, similarly to Figure 2. The darker shades show three profiles with the default thermal Planck opacity presented in the results section, $\kappa_{\mathrm{P,therm}}=7.5$ cm2 g-1, and solar opacities such that the ratio between the two, $\gamma$ is 0.3 (green), 3 (blue), and 30 (purple). The lighter shades show the same three choices of $\gamma$, but with both opacities decreased by a factor of 10. The mass loss rates for the same value of $\gamma$ are linked by solid lines in the right panel. We find that lower opacities move the $\tau=1$ surfaces inward for the same $\gamma$. This moves the transition point from $T_{\mathrm{in}}$ to $T_{\mathrm{out}}$ inward. If $\gamma>1$, then $T_{\mathrm{out}}>T_{\mathrm{in}}$, and the larger region of the atmosphere at higher temperatures leads to an increase in the mass loss rate. Conversely, if $\gamma<1$, then the mass loss rate decreases. This increase can be up to a factor of 5 for $\gamma=30$, but since $\gamma$ still determines the relative temperatures, the mass loss rates scale with $\gamma$ in a similar way. We therefore conclude that while the thermal structure, and therefore the mass loss rate, is sensitive to the choice of absolute opacities, such variation is second-order compared to the value of $\gamma$, which remains the main control on mass loss rate.
@@ Acknowledgements
This research has been supported by NASA’s Exoplanet Research Program (XRP) under grant number 80NSSC21K0392. WM is also supported by a UCLA Dissertation Year Fellowship. JEO is supported by a Royal Society University Research Fellowship. JEO and MS have received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement no. 853022, PEVAP).
@@ Appendix B Tabulated mass loss rates
Table 1: The first entries in the tabulated mass loss rates tables, for a planet with $M_{\mathrm{p}}=5M_{\oplus}$ and $T_{\mathrm{eq}}=1000$ K. Each row corresponds to a unique aiolos run. The first column is the ratio of the thermal and stellar opacities, $\gamma$, used in the run, while the second and third columns are the specific Planck opacity to thermal, $\kappa_{\mathrm{P,therm}}$, and stellar, $\kappa_{\mathrm{P,\odot}}$ radiation, both in cm2 g-1. The fourth column is the transit radius, $R_{\mathrm{trans}}$, calculated according to Eq. 8, in cm. The fifth column is the resultant mass loss rate, $\dot{M}$, in g s-1. These tables can be interpolated to calculate the bolometrically-driven mass loss rate as functions of opacity ratio, observed radius, core mass, and equilibrium temperature. A machine-readable version of this table and those for other combinations of planet mass and equilibrium temperature are available as Supplementary Data.

### v2

@@ 4.1 Sensitivity of mass loss rates to opacity
Figure 2: Temperature (top), density (middle), and velocity (bottom) profiles as functions of radius, in core radii $R_{\mathrm{c}}$, and corresponding mass loss rates (right) for seven values of the opacity ratio ranging from 0.03 to 30 (different colors from yellow to purple). All profiles have a fixed radiative-convective boundary location $R_{\mathrm{rcb}}=2R_{\mathrm{c}}=1.91\times 10^{9}$ cm, planet mass $M_{\mathrm{p}}=5M_{\oplus}$, atmospheric mass fraction $f=0.025$, and incident flux equivalent to an equilibrium temperature $T_{\mathrm{eq}}=1000\,$K (shown by the gray horizontal line on the top panel). T-shaped markers represent the transit radius, while squares represent the sonic radius. Shown in red dashed lines are the equivalent profiles and mass loss rates for an isothermal radiative region at the analytic interior temperature, $T_{\mathrm{in}}=T_{\mathrm{eq}}/2^{1/4}$. Increasing the opacity ratio leads to different temperature and density profiles and therefore to mass loss rates that vary two orders of magnitude over plausible values. The mass loss rate for the $\gamma=1$ profile are similar to those for an isothermal outer atmosphere at $T_{\mathrm{in}}$.
@@
In Fig. 2, we compare the thermal (top left), density (middle left), and velocity (bottom left) profiles and mass loss rates (right) as functions of the ratio of the opacity to incident visible and thermal radiation, $\gamma$. Each color is a run with a value of this ratio ranging from 0.3 to 30, with $\gamma$ increasing as the colors get darker. All seven runs have a fixed radiative-convective boundary radius, $R_{\mathrm{rcb}}=2R_{\mathrm{c}}=1.91\times 10^{9}$ cm, a fixed initial atmospheric mass of $0.025M_{\mathrm{c}}$, a planet mass $M_{\mathrm{p}}=5M_{\oplus}$, and a fixed incident flux equivalent to an equilibrium temperature of 1000 K, which is shown by the gray horizontal line. For each $\gamma$ value, the solid line represents the output of the aiolos code, while the dotted line represents the semi-analytic extrapolation into the optically thick region described in Section 3.6. T-shaped markers represent the transit radius, $R_{\mathrm{trans}}$ (Eq. 2), while squares represent the sonic radius, $R_{\mathrm{s}}$ (Eq. 11). The red dashed lines and red diamond represent results using an isothermal profile: a hydrostatic density profile with the same radiative-convective boundary radius and atmospheric mass, a constant temperature equal to $T_{\mathrm{in}}=T_{\mathrm{eq}}/2^{1/4}$, a Parker-type velocity profile following Cranmer (2004), and the corresponding mass loss rate following Eq. 10.
@@
In general, the numerical thermal profiles reproduce the characteristics described in Section 2, thus looking broadly similar to the schematic in Fig. 1. As the opacity ratio $\gamma$ increases, the outer region, optically thin to incident radiation, heats up, with temperatures increasing by a factor of 7 over the studied 1000-fold opacity ratio increase. The outer temperatures range from well below the equilibrium temperature to well above it, and are well-predicted by the $T_{\mathrm{out}}=T_{\mathrm{eq}}(\gamma/4)^{1/4}$ relation of Guillot (2010) and Schulik & Booth (2023). Meanwhile the inner region, optically thick to outgoing radiation, reaches lower temperatures as $\gamma$ is increased, decreasing by 100 K over the studied range. In all cases, the inner, optically thick radiative region is a few hundred kelvins cooler than the equilibrium temperature. This temperature is also lower by $\sim 70-200$ K than the analytic lower limit of $T_{\mathrm{in}}$ discussed in Schulik & Booth (2023). The minimum temperature also increases monotonically with $\gamma$, rather than being highest at $\gamma\sim 1$. This is because the inner temperature approximation of Parmentier & Guillot (2014) was calculated assuming plane-parallel geometry. However, in small planets the scale height is a non-negligible fraction of the planet radius. Therefore, in our simulations the different surface areas of the absorption and emission surfaces lead to a decrease in temperature as $\kappa_{P,\odot}$ increases, since the same radiation is absorbed over a progressively larger area.
@@ 5.2 Comparison to direct observations
A more direct test of the thermal structure of sub-Neptunes and its effects on planet evolution could come from atmospheric escape measurements. Signatures of escaping atmospheres have been observed directly for sub-Neptunes. Detections consistent with atmospheric escape for small exoplanets have been made using hydrogen Ly $\alpha$ 121 nm in the UV (e.g. Kulow et al., 2014; Ehrenreich et al., 2015) and He 1083 nm in the NIR (e.g. Mansfield et al., 2018) lines (see Dos Santos (2023) for a recent review). If these planets are undergoing bolometrically-driven escape, measurements of the mass loss rate will constrain the temperature of the upper atmosphere, and therefore the atmospheric opacity ratio. However, the mass loss rate is difficult to constrain precisely from these measurements, due to degeneracies between escape velocity and ionization by the stellar wind (Owen et al., 2023; Schreyer et al., 2024). But using existing models, these measurements have been fit by isothermal mass loss rate estimates for sub-Neptunes as low as $10^{9}$ g s-1 (Ehrenreich et al., 2015; Mansfield et al., 2018). This measured rate is within the range of escape rates we predict on billion year timescales for some planet-$\gamma$ combinations, indicating that directly testing upper atmospheric structure with escape observations is within reach of existing telescopic capabilities. We facilitate comparisons to our results by tabulating the mass loss rates we obtain with aiolos as functions of planet mass, equilibrium temperature, value of $\gamma$, and transit radius and including them as machine-readable Supplementary Data. We describe these tables in Appendix B.
@@ 5.3 Opacity variations in space and time
Figure 9: Temperature (top), density (middle), and velocity (bottom) profiles as functions of radius, in core radii $R_{\mathrm{c}}$, and corresponding mass loss rates (right) for six choices of Planck opacities. In the darker colors are three profiles with the default thermal Planck opacity presented in Section 4, $\kappa_{\mathrm{P,therm}}=7.5$ cm2 g-1, and solar opacities such that the ratio between the two, $\gamma$ is 0.3 (green), 3 (blue), and 30 (purple). These profiles were previously shown in Fig. 2. The lighter shades of each color correspond to runs with the same $\gamma$ values, but all opacities lowered by a factor of 10, such that $\kappa_{\mathrm{P,therm}}=0.75$ cm2 g-1. In the right panel, the mass loss rates for each particular value of $\gamma$ are linked by solid lines. All opacities listed in text on the figure are in cm2 g-1. All profiles have a fixed radiative-convective boundary location $R_{\mathrm{rcb}}=2R_{\mathrm{c}}=1.91\times 10^{9}$ cm, atmospheric mass fraction $f=0.025$, and incident flux equivalent to an equilibrium temperature $T_{\mathrm{eq}}=1000\,$K (shown by the gray horizontal line on the top panel). Shown in red dashed lines are the equivalent profiles and mass loss rates for an isothermal radiative region at the analytic interior temperature, $T_{\mathrm{in}}=T_{\mathrm{eq}}/2^{1/4}$. As in Fig. 2, T-shaped markers represent the transit radius, while squares represent the sonic radius. Decreasing the opacities moves the $\tau=1$ surfaces inward, and therefore the transitions between $T_{\mathrm{in}}$ and $T_{\mathrm{out}}$ move inward for each $\gamma$. For $\gamma>1$, where $T_{\mathrm{out}}>T_{\mathrm{in}}$, this increases the mass loss rate, up to a factor of 5 for the $\gamma=30$ case shown, while it lowers the mass loss rate for $\gamma<1$. The value of $\gamma$ still sets the outer temperatures and therefore remains the main control on the mass loss rate.
@@
In order to ascertain that our results are independent to first-order of our exact choice of the thermal Planck opacity $\kappa_{\mathrm{P,therm}}$, we run simulations with a lower value, $\kappa_{\mathrm{P,therm}}=0.75$ cm2 g-1, shown in Figure 9. Here we show the temperature, density, and velocity profiles as functions of radius, as well as the mass loss rate, similarly to Figure 2. The darker shades show three profiles with the default thermal Planck opacity presented in the results section, $\kappa_{\mathrm{P,therm}}=7.5$ cm2 g-1, and solar opacities such that the ratio between the two, $\gamma$ is 0.3 (green), 3 (blue), and 30 (purple). The lighter shades show the same three choices of $\gamma$, but with both opacities decreased by a factor of 10. The mass loss rates for the same value of $\gamma$ are linked by solid lines in the right panel. We find that lower opacities move the $\tau=1$ surfaces inward for the same $\gamma$. This moves the transition point from $T_{\mathrm{in}}$ to $T_{\mathrm{out}}$ inward. If $\gamma>1$, then $T_{\mathrm{out}}>T_{\mathrm{in}}$, and the larger region of the atmosphere at higher temperatures leads to an increase in the mass loss rate. Conversely, if $\gamma<1$, then the mass loss rate decreases. This increase can be up to a factor of 5 for $\gamma=30$, but since $\gamma$ still determines the relative temperatures, the mass loss rates scale with $\gamma$ in a similar way. We therefore conclude that while the thermal structure, and therefore the mass loss rate, is sensitive to the choice of absolute opacities, such variation is second-order compared to the value of $\gamma$, which remains the main control on mass loss rate.
@@ Acknowledgements
This research has been supported by NASA’s Exoplanet Research Program (XRP) under grant number 80NSSC21K0392. WM has also been supported by a UCLA Dissertation Year Fellowship. JEO is supported by a Royal Society University Research Fellowship. JEO and MS have received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement no. 853022, PEVAP). We thank the anonymous reviewer for their comments, which improved the clarity of the manuscript.
@@ Appendix B Tabulated mass loss rates
Table 1: Tabulated Mass Loss Rates